OpenStudy (anonymous):

Write the formula given the following arithmetic sequences. 1.001, 2.012, 3.023, 4.034, ... tn = 1.001 + 1.011(n – 1) tn = 1 + 1.001(n – 1) tn = 1.001 + 1.010(n – 1) tn = 1.001 – 1.011(n – 1)

OpenStudy (anonymous):

a+(n-1)d = 1.001+1.011(n-1)

OpenStudy (cwrw238):

a = 1 .001 d = 2.012 - 1 001 = 1.011 tn = 1.001 + 1.011(n - 1)